OpenStudy (anonymous):

Evaluate the integral \[\int\limits_{?}^{?} \frac{dx}{1-\sin x \cos x} \] I've been trying this for a while, any help?

OpenStudy (anonymous):

I tried doing \[\cos x = \sqrt{1-\sin^2 x}\] And then \[u=\sin(x)\] But it didn't really help...

OpenStudy (zarkon):

you might want to try the Weierstrass substitution http://planetmath.org/encyclopedia/WeierstrassSubstitutionFormulas.html

OpenStudy (zarkon):

I haven't tried it yet so I don't know how will it will work on this problem

hero (hero):

\[\int\limits \frac{1}{1-\sin x \cos x} dx\]

OpenStudy (zarkon):

that solution uses the Weierstrass substitution ;)

OpenStudy (anonymous):

Thanks... I actually posted this on behalf of a friend, I have no clue why he needs to do this in highschool O_O

hero (hero):

AP Calculus?